# Real-IAD Variety: Pushing Industrial Anomaly Detection Dataset to a Modern Era

### v1

@@ 4.3 Benchmark on Zero-Shot and Few-Shot AD
Experimental Setting. Zero-Shot Anomaly Detection (ZSAD) requires models trained on auxiliary datasets to identify anomalies in novel object categories without class-specific training data. This setting is particularly valuable because obtaining comprehensive coverage of all object classes and anomaly types is impractical in real-world scenarios. ZSAD aims to enhance model robustness and adaptability by facilitating knowledge transfer from known to unknown categories, thereby enabling effective detection of novel anomalies across diverse and dynamic environments. To further evaluate model adaptability in open-world scenarios, we introduce a Few-Shot Anomaly Detection (FSAD) setting alongside ZSAD evaluation. In FSAD, a limited number of normal images ($k=\{1,2\}$) from the target class serve as visual prompts during inference, with their detection results integrated with ZSAD outputs to enhance final anomaly detection performance.
@@
Competitive Methods. We evaluate AnomalyCLIP [7], AdaCLIP [38], VCP-CLIP [39], and AdaptCLIP [8] on Real-IAD Variety in zero-shot manner, conducting comprehensive comparisons with Real-IAD [17]. Given that Real-IAD Variety comprises 160 novel categories, it serves as an exceptionally challenging benchmark for assessing ZSAD model robustness and cross-category generalization in complex real-world scenarios. Notably, AdaptCLIP [8] supports both zero- and few-shot generalization across domains; therefore, we employ it to evaluate FSAD performance on Real-IAD Variety. Results and Analysis under ZSAD and FSAD. Table 4 presents ZSAD and FSAD results on Real-IAD and Real-IAD Variety, with performance trends visualized in Figure 4. Several key observations emerge:
@@
ZSAD models demonstrate remarkable stability and robustness, while traditional MUAD methods exhibit significant performance degradation as category complexity increases. This fundamental difference is quantitatively highlighted by low I-AUROC delta values (e.g., 2.3 for AnomalyCLIP, 1.6 for AdaptCLIP) across increasing category counts (30 to 160). This suggests that VLM-based ZSAD approaches, which leverage external knowledge from vast pre-training data rather than modeling intrinsic training set distributions, are inherently less sensitive to data distribution shifts induced by category scale-up.
FSAD yields substantial performance gains using minimal class-specific prompt information. Specifically, using only two normal images (2-Shot), AdaptCLIP’s performance on Real-IAD Variety (160 classes) improves significantly, with I-AUROC rising from 73.0% (ZSAD) to 81.4% (2-Shot). This improvement confirms the high value of incorporating even minimal normal visual prompts to rapidly refine generic VLM knowledge for precise, target-specific anomaly discrimination. The consistent improvement from 1-Shot to 2-Shot further validates the effectiveness of this adaptation mechanism.
Few-shot AdaptCLIP (2-Shot) achieves the best overall performance across all benchmarks. On the most challenging 160-class Real-IAD Variety, 2-Shot AdaptCLIP reaches I-AUROC of 81.4%, which is highly competitive with the best MUAD model (Dinomaly+: 81.6%) that utilizes full normal training data. Crucially, delta analysis reveals that while few-shot adaptation boosts absolute performance, it incurs slightly higher delta for pixel-level metrics, suggesting that local adaptation mechanisms, while powerful, exhibit greater sensitivity to category complexity than global ZSAD paradigms.
The superior performance and inherent stability of VLM-based ZSAD/FSAD methods strongly advocate for multimodal knowledge integration. To further unlock the potential of these models, particularly in the complex cross-category setting of Real-IAD Variety, future research should focus on developing enhanced methods that effectively utilize detailed textual and semantic information, bridging the gap between high-level semantic understanding and fine-grained localized visual anomaly detection.
@@ 4.4 Benchmark on Multi-View AD
Experimental Setting. Multi-View Anomaly Detection (MVAD) originates from high-precision demands in industrial quality inspection. It aims to achieve detection capabilities unattainable from single-view images by integrating multiple perspective images of the same sample, thereby identifying defects that are invisible from certain viewpoints but visible from others. Additionally, this setting leverages multi-view information and consistency constraints to further enhance model performance.
Table 5: MVAD [35] performance on Real-IAD Variety with increasing categories.
Datasets # Classes I-AUROC P-AUROC P-AUPR
Real-IAD 30 86.6 97.9 30.3
Real-IAD Variety S1 30 85.5 91.9 34.7
Real-IAD Variety S2 60 82.9 91.6 29.9
Real-IAD Variety S3 100 81.8 91.5 26.8
Real-IAD Variety 160 78.1 90.4 22.3
@@ 5 Future Work
Multimodal Anomaly Detection with Large-Scale Diverse Data. The introduction of Real-IAD Variety opens several promising research directions for advancing anomaly detection. First, the development of robust generalized anomaly detection models represents a critical avenue. Future research should explore training strategies that effectively integrate these textual insights with visual features, e.g., MMAD [45], thereby enhancing the robustness and adaptability of anomaly detection systems. Second, addressing logical anomalies in image-based anomaly detection remains an important challenge. Current methods often struggle with logical anomalies, where visual representations alone fail to capture complete contextual relationships among objects or components. Our preliminary ZSAD and FSAD experiments have revealed performance limitations in this regard. To address these constraints, we propose enabling text encoder fine-tuning to enhance model adaptability and contextual reasoning capabilities. Future research should focus on integrating textual descriptions into existing anomaly detection frameworks, enabling models to reason about logical relationships and contextual cues beyond purely visual patterns.
@@ 6 Conclusion
Industrial Anomaly Detection (IAD) is undergoing a fundamental transformation with the emergence of large-scale models capable of operating in unified and zero-/few-shot settings. This work makes a substantial contribution to the IAD dataset landscape by introducing Real-IAD Variety, a large-scale benchmark that not only expands the scope of available data but also establishes rigorous evaluation protocols for assessing state-of-the-art (SOTA) IAD methodologies. Real-IAD Variety, characterized by its unprecedented scale and diversity, redefines the benchmark standards for IAD datasets. It encompasses 160 categories spanning 28 industries, 24 material types, and 22 color variations with 198,960 images, effectively overcoming the limitations of previous datasets that were constrained by narrow category ranges and limited scenario representation.
Our comprehensive experimental analysis on Real-IAD Variety reveals several critical findings: (1) Multi-class unsupervised anomaly detection (MUAD) methods experience significant performance degradation as category count increases, with performance declining by $10–30\%$ when scaling from 30 to 160 categories; (2) Zero-shot and few-shot approaches demonstrate remarkable resilience to category scale-up, with minimal performance variation across different category counts, suggesting their superior effectiveness in handling diverse industrial scenarios; (3) Vision-language models (VLMs) leveraging external pre-trained knowledge exhibit fundamentally different scalability characteristics compared to traditional unsupervised methods, highlighting the importance of multimodal learning for large-scale IAD applications. This research underscores the pivotal role of dataset diversity and scale in advancing IAD model development. Real-IAD Variety and its associated benchmarks provide a solid foundation for future research and innovation, particularly in multi-class, multi-view and zero-/few-shot anomaly detection paradigms. We anticipate that this benchmark will facilitate the development of models with enhanced generalization capabilities, ultimately improving the reliability and efficiency of anomaly detection systems in real-world industrial applications.

### v2

@@ 4.3 Benchmark on Zero-Shot and Few-Shot AD
Experimental Setting. Zero-Shot Anomaly Detection (ZSAD) requires models trained on auxiliary datasets to identify anomalies in novel object categories without class-specific training data. This setting is particularly valuable because obtaining comprehensive coverage of all object classes and anomaly types is impractical in real-world scenarios. ZSAD aims to enhance model robustness and adaptability by facilitating knowledge transfer from known to unknown categories, thereby enabling effective detection of novel anomalies across diverse and dynamic environments. To further evaluate model adaptability in open-world scenarios, we introduce a Few-Shot Anomaly Detection (FSAD) setting alongside ZSAD evaluation. In FSAD, a limited number of normal images ($k\in\{1,4\}$) from the target class serve as visual prompts during inference, with their detection results integrated with ZSAD outputs to enhance final anomaly detection performance. To ensure statistical significance, performance for all FSAD models is averaged across three runs with different random seeds, reporting both mean and standard deviation.
@@
Competitive Methods. We evaluate AnomalyCLIP [49], AdaCLIP [2], VCP-CLIP [31], and AdaptCLIP [10] on Real-IAD Variety in a zero-shot manner, conducting comprehensive comparisons with Real-IAD [39]. Given that Real-IAD Variety comprises 160 novel categories, it serves as an exceptionally challenging benchmark for assessing ZSAD model robustness and cross-category generalization in complex real-world scenarios. Notably, AdaptCLIP [10] supports both zero- and few-shot generalization across domains. Therefore, we employ it to evaluate FSAD performance on Real-IAD Variety. In addition, we also evaluate the one-shot pure-vision MetaUAS [12].
Results and Analysis under ZSAD and FSAD. Table 5 presents ZSAD and FSAD results on Real-IAD and Real-IAD Variety, with performance trends visualized in Fig. 6.
@@
Some key observations are as follows: (1) ZSAD models demonstrate remarkable stability and robustness, while traditional MUAD methods exhibit significant performance degradation as category complexity increases. This fundamental difference is quantitatively highlighted by low I-ROC $\Delta$ values (e.g., 0.9 for AnomalyCLIP, 0.4 for AdaptCLIP) across increasing category counts (30 to 160). This suggests that VLM-based ZSAD approaches, which leverage external knowledge from vast pre-training data rather than modeling intrinsic training set distributions, are inherently less sensitive to data distribution shifts induced by category scale-up. (2) FSAD yields substantial performance gains by leveraging class-specific prompt information. Specifically, using only 4 normal images (4-Shot), AdaptCLIP’s performance on Real-IAD Variety (160 classes) improves significantly, with I-AUROC rising from 72.9% (ZSAD) to 88.1% (4-Shot). This improvement confirms the high value of incorporating even few-shot normal visual prompts to rapidly refine generic VLM knowledge for precise, target-specific anomaly discrimination. The consistent improvement from 1-Shot to 4-Shot further validates the effectiveness of this adaptation mechanism. (3) Few-shot AdaptCLIP (4-Shot) achieves the best overall performance across all benchmarks. On the most challenging 160-class Real-IAD Variety, 4-Shot AdaptCLIP reaches I-ROC of 88.1%, which is highly competitive with the best multi-class Dinomaly [16] (85.4%) that utilizes full normal training data. Crucially, delta analysis reveals that while few-shot adaptation boosts absolute performance, it incurs slightly higher delta for pixel-level metrics, suggesting that local adaptation mechanisms, while powerful, exhibit greater sensitivity to category complexity than global ZSAD paradigms.
ZSAD and FSAD leverage semantic anchoring via textual prompts or visual representation with normal images prompts. Unlike MUAD, these methods utilize a pre-aligned latent space where linguistic concepts of anomalies (e.g., “damaged”, “defect”) provide a unified representation of common anomalous (normal visual) features. Because they rely on frozen VLMs and semantic reasoning rather than statistical learning on specific datasets, they require no category-specific fine-tuning. Consequently, their generalization capability is fundamentally decoupled from the number of test categories. MUAD models may suffer from feature “pollution” or capacity constraints as more classes are added. In contrast, ZSAD models identify defects based on pre-trained semantic alignment rather than statistical novelty relative to a specific training set. To further unlock the potential of these models, particularly in the complex cross-category setting of Real-IAD Variety, future research should focus on developing enhanced methods that effectively utilize detailed textual and semantic information, bridging the gap between high-level semantic understanding and fine-grained localized visual anomaly detection.
@@ 4.4 Benchmark on Multi-View AD
Experimental Setting. Multi-View Anomaly Detection (MVAD) originates from high-precision demands in industrial quality inspection. It aims to achieve detection capabilities unattainable from single-view images by integrating multiple perspective images of the same sample, thereby detecting defects that may be occluded or invisible in certain viewpoints but remain discernible in others. Additionally, this setting leverages multi-view information and consistency constraints to further enhance model performance.
Table 6: MVAD [18] performance comparison of Real-IAD Variety with increasing categories.
Model/Dataset I-ROC I-PR I-F1max P-ROC P-PR P-F1max P-PRO
Real-IAD-30c 86.6 84.8 77.2 97.9 30.3 36.8 91.2
Real-IAD Variety S1 85.4 97.0 94.8 90.7 36.8 39.7 72.4
Real-IAD Variety S2 81.4 96.2 94.1 90.3 31.8 36.2 71.0
Real-IAD Variety S3 80.3 95.8 94.0 90.5 28.8 33.9 70.5
Real-IAD Variety 77.1 95.1 93.6 89.5 24.5 30.2 67.7
@@
4.5 Potential for Foundational AD Models
Table 7: Few-shot anomaly detection results on MVTec and VisA. We compare the ADPretrain model (FeatureNorm) pre-trained on Real-IAD versus our Real-IAD Variety.
Test Set Train Set Shot I-ROC I-PR I-F1max P-ROC P-PR P-F1max P-PRO
MVTec Real-IAD 1 87.9 93.8 90.2 96.7 48.1 53.4 91.4
2 95.6 97.0 95.3 97.3 50.2 55.2 92.3
4 97.4 98.3 96.6 97.4 50.6 55.6 92.5
8 97.8 98.6 96.8 97.5 51.5 56.0 92.8
Real-IAD Variety 1 90.7 93.8 93.1 97.3 54.1 56.6 92.0
2 96.0 97.5 95.8 97.7 56.1 58.1 92.9
4 97.7 99.0 97.2 98.0 56.9 58.6 93.6
8 98.1 99.1 97.5 98.1 57.4 59.1 93.7
ViSA Real-IAD 1 61.4 65.7 73.2 96.4 19.2 28.0 81.0
2 68.8 75.1 74.1 97.0 25.1 33.5 83.7
4 75.2 81.5 77.2 97.6 28.8 37.3 85.5
8 82.6 84.1 80.4 98.0 31.7 40.9 87.1
Real-IAD Variety 1 58.6 62.4 73.1 96.8 20.8 30.7 83.1
2 70.4 74.6 75.1 97.6 30.0 37.2 86.5
4 88.1 90.0 84.8 98.0 34.4 40.5 88.5
8 93.4 94.4 89.6 98.3 36.1 42.1 89.7
Experimental Setting. To demonstrate the generalization of foundational AD models pre-trained large-scale Real-IAD Variety dataset. We employ ADPretrain [40] framework (DINOv2-base backbone) and compare the effects of pre-training on the original Real-IAD and our proposed Real-IAD Variety. Then, we evaluate few-shot (1, 2, 4, 8-shot) generalization on the independent MVTec and VisA benchmarks.
Few-shot Generalization. As shown in Table 7, the few-shot performance of ADPretrain model pre-trained on Real-IAD Variety consistently outperforms the baseline across all settings. Notably, the performance gains are significantly more pronounced on the VisA dataset, which is widely recognized as a more challenging benchmark due to its complex object structures and subtle defects. For instance, in the 4-shot setting on VisA, pre-training on Real-IAD Variety improves I-ROC from 75.2% to 88.1% (+12.9%). We observe a slight fluctuation in the 1-shot ViSA setting, which may be attributed to the high sensitivity of extremely low-shot sampling in certain complex categories. These results empirically demonstrate that the increased category diversity and broader industrial scope of Real-IAD Variety are crucial for enhancing generalization on foundational AD models. By providing a superior foundation for feature alignment, our dataset helps foundational models learn robust visual primitives that generalize far better to challenging unseen domains.
@@ 5 Future Work
Multimodal Anomaly Detection with Large-Scale Diverse Data. The introduction of Real-IAD Variety opens several promising research directions for advancing anomaly detection. First, the development of robust generalized anomaly detection models represents a critical avenue. Future research should explore training strategies that effectively integrate these textual insights with visual features, e.g., MMAD [23], thereby enhancing the robustness and adaptability of anomaly detection systems. Second, addressing logical anomalies in image-based anomaly detection remains an important challenge. Current methods often struggle with logical anomalies, where anomalies are defined by misplaced components or incorrect counts rather than distinctive surface textures. Our ZSAD/FSAD results suggest that while VLMs excel at semantic alignment, bridging the gap between high-level reasoning and fine-grained spatial logic remains a frontier for foundational IAD research. Our preliminary ZSAD and FSAD experiments have revealed performance limitations in this regard. To address these constraints, we propose enabling text encoder fine-tuning to enhance model adaptability and contextual reasoning capabilities. Future research should focus on integrating textual descriptions into existing anomaly detection frameworks, enabling models to reason about logical relationships and contextual cues beyond purely visual patterns.
@@ 6 Conclusion
Industrial Anomaly Detection (IAD) is undergoing a fundamental transformation with the emergence of large-scale models capable of operating in unified and zero-/few-shot settings. This work makes a substantial contribution to the IAD dataset landscape by introducing Real-IAD Variety, a large-scale benchmark that not only expands the scope of available data but also establishes rigorous evaluation protocols for assessing state-of-the-art (SOTA) IAD methodologies. Real-IAD Variety, characterized by its unprecedented scale and diversity, redefines the benchmark standards for IAD datasets. It encompasses 160 categories spanning 28 industries, 24 material types, 22 color variations and 27 defect types with 198,950 images, effectively overcoming the limitations of previous datasets that were constrained by narrow category ranges and limited scenario representation.
Our comprehensive experimental analysis on Real-IAD Variety reveals several critical findings: (1) Most multi-class unsupervised anomaly detection methods experience significant performance degradation as the category count increases, with I-ROC typically declining by 10%–20% when scaling from 30 to 160 categories; (2) Zero-shot and few-shot approaches demonstrate remarkable resilience to category scale-up, exhibiting minimal performance fluctuation. This suggests their superior scalability and effectiveness in handling diverse industrial scenarios; (3) Vision-language models (VLMs) leveraging external pre-trained knowledge exhibit fundamentally different scalability characteristics compared to traditional unsupervised methods, highlighting the pivotal role of multimodal learning in the next generation of IAD systems. Real-IAD Variety and its associated benchmarks provide a solid foundation for future research and innovation, particularly in multi-class, multi-view and zero-/few-shot anomaly detection paradigms. We anticipate that this benchmark will facilitate the development of models with enhanced generalization capabilities, ultimately improving the reliability and efficiency of anomaly detection systems in real-world industrial applications.